### v1

# Attention Isn’t All You Need for Emotion Recognition: Domain Features Outperform Transformers on the EAV Dataset

Abstract
We present a systematic study of multimodal emotion recognition using the EAV dataset, investigating whether complex attention mechanisms improve performance on small datasets. We implement three model categories: baseline transformers (M1), novel factorized attention mechanisms (M2), and improved CNN baselines (M3). Our experiments show that sophisticated attention mechanisms consistently underperform on small datasets. M2 models achieved 5 to 13 percentage points below baselines due to overfitting and destruction of pretrained features. In contrast, simple domain-appropriate modifications proved effective: adding delta MFCCs to the audio CNN improved accuracy from 61.9% to 65.56% (+3.66pp), while frequency-domain features for EEG achieved 67.62% (+7.62pp over the paper baseline). Our vision transformer baseline (M1) reached 75.30%, exceeding the paper’s ViViT result (74.5%) through domain-specific pretraining, and vision delta features achieved 72.68% (+1.28pp over the paper CNN). These findings demonstrate that for small-scale emotion recognition, domain knowledge and proper implementation outperform architectural complexity.
Attention Isn’t All You Need for Emotion Recognition:
Domain Features Outperform Transformers on the EAV Dataset
Anmol Guragain Universidad Politécnica de Madrid E.T.S. Ingenieros de Telecomunicación anmol.g@upm.es
1 Introduction
Emotion recognition from physiological and behavioral signals is a core challenge in affective computing, with applications in human-computer interaction, mental health monitoring, and assistive technologies. The EAV dataset Kim et al. (2024) provides synchronized EEG, Audio, and Video recordings from 42 participants engaged in conversational interactions, offering a unique resource for multimodal emotion research. However, with only approximately 280 training samples per subject after data splitting, this dataset presents significant challenges for deep learning approaches.
This limited data availability motivates our central research question: Do complex attention mechanisms improve emotion recognition performance on small datasets, or do simpler approaches with domain-appropriate features perform better?
To answer this question, we systematically evaluate three categories of models. First, we implement baseline transformer architectures (M1) to establish reference performance. Second, we develop novel factorized attention mechanisms (M2) tailored to each modality’s unique structure: spatial-temporal-asymmetry attention for EEG, temporal-frequency dual attention for audio, and ViViT-style space-time attention for video. Third, we explore minimal CNN improvements (M3) focusing on bug fixes and domain-appropriate feature engineering.
Our study makes three contributions: (1) we show that factorized attention mechanisms consistently fail on small datasets, underperforming baselines by 5 to 13 percentage points; (2) we demonstrate that simple, domain-appropriate modifications (delta MFCCs for audio and band power features for EEG) achieve state-of-the-art results on this dataset; and (3) we provide analysis of failure modes and success factors for emotion recognition on limited data.
2 Related Work
2.1 Emotion Recognition Datasets
Traditional EEG datasets such as DEAP Koelstra et al. (2012) and MAHNOB-HCI employ passive elicitation paradigms where participants watch emotional stimuli. Audio-visual datasets like IEMOCAP Busso et al. (2008) capture conversational emotions but lack neurophysiological signals. The EAV dataset bridges this gap by providing synchronized multimodal recordings during active conversation, though its relatively small scale (42 participants) constrains the complexity of applicable models.
2.2 Transformer Architectures for Emotion Recognition
The Audio Spectrogram Transformer (AST) Gong et al. (2021) treats spectrograms as images, applying Vision Transformer architecture with strong performance on large-scale audio datasets. Video Vision Transformer (ViViT) Arnab et al. (2021) extends this approach to video through factorized space-time attention, achieving 22 computational efficiency over full attention. EEGformer Hussain and Park (2022) adapts transformers for EEG signals. These architectures achieve impressive results on large datasets, but their effectiveness on small-scale emotion recognition remains unclear.
2.3 Frontal Asymmetry in Emotion
Davidson’s frontal asymmetry model Davidson (1992) establishes that left frontal cortex activation (measured at F3) correlates with approach motivation and positive emotions, while right frontal activation (F4) correlates with withdrawal motivation and negative emotions. This neurophysiological insight motivates our M2 EEG architecture and M3 band power features.
2.4 Feature Engineering for Emotion Recognition
Classical approaches to emotion recognition rely on hand-crafted features. For EEG, band power and differential entropy features computed from frequency bands (delta, theta, alpha, beta, gamma) have proven effective on the SEED dataset Zheng and Lu (2015). For audio, Mel-frequency cepstral coefficients (MFCCs) and their temporal derivatives (delta and delta-delta) capture both spectral content and dynamics Schuller and others (2013). Squeeze-and-Excitation networks Hu et al. (2018) introduced channel attention for CNNs, learning to weight feature map importance through a bottleneck architecture.
3 Dataset
The EAV dataset contains recordings from 42 participants, each providing 200 interaction samples across five emotion classes: Neutral, Anger, Happiness, Sadness, and Calmness. EEG signals are recorded using a 30-channel BrainAmp system at 500 Hz, audio at 16 kHz, and video at 30 fps.
Following standard preprocessing, EEG signals are downsampled to 100 Hz and bandpass filtered (0.5 to 45 Hz), yielding segments of shape (30 channels, 500 samples). Audio clips of 5 seconds are extracted, and 25 video frames are sampled per segment. The data is split 70/30 for training and testing per subject, resulting in approximately 280 training and 120 test samples per subject per modality.
This limited training data (280 samples with 15,000 EEG features, 80,000 audio samples, or 252242243 video pixels) fundamentally constrains model complexity and motivates our investigation of simpler approaches.
4 Methods
4.1 M1: Baseline Transformer Architectures
Our baseline models replicate the paper’s transformer approach using pretrained architectures from Hugging Face.
EEG Transformer.
We implement a custom architecture with convolutional feature extraction followed by transformer encoding. Two 1D convolutional layers (kernel size 11, padding 5) transform the 30-channel input into 60-dimensional features while preserving temporal resolution. Six transformer encoder layers with 4 attention heads and hidden dimension 60 model temporal dependencies. Mean pooling followed by a two-layer classifier produces emotion predictions. This architecture applies standard self-attention across time, treating channels as input features rather than a separate dimension.
Audio Spectrogram Transformer.
We fine-tune the pretrained AST model (MIT/ast-finetuned-audioset) with a custom 5-class classifier head. The model converts audio waveforms to mel spectrograms (1024128), extracts 1616 patches yielding approximately 1212 tokens, and applies 12 transformer layers with 768-dimensional embeddings. Training proceeds in two stages: 10 epochs with frozen backbone (learning rate 5e-4), followed by 15 epochs with unfrozen backbone (learning rate 5e-6).
Vision Transformer.
We employ a ViT model pretrained on facial emotion recognition (dima806/facial_emotions_image_detection). Each video is processed as 25 independent frames, with per-frame predictions averaged for the final classification. The same two-stage training strategy is applied: 10 epochs frozen, then 5 epochs unfrozen. While this approach ignores temporal dynamics across frames, the facial emotion pretraining provides strong baseline performance.
4.2 M2: Factorized Attention Mechanisms
Our main architectural contribution explores whether factorized attention, tailored to each modality’s structure, can improve upon baselines. Figure 1 illustrates the three M2 architectures.
4.2.1 M2-EEG: Tri-Stream Transformer
EEG signals contain information along three distinct axes. The spatial dimension captures relationships between brain regions (channels), the temporal dimension captures signal dynamics, and the hemispheric asymmetry dimension captures left-right differences relevant for emotional valence.
Our architecture first applies convolutional feature extraction, then branches into three parallel attention streams. The spatial stream transposes features to (time, channels, features) and applies attention across channels, learning which brain regions interact at each time point. The temporal stream applies attention across time points for each channel, capturing event-related dynamics. The asymmetry stream computes difference signals between six hemisphere pairs (Fp1-Fp2, F3-F4, F7-F8, C3-C4, P3-P4, O1-O2) based on the 10-20 electrode system, then applies attention across these pairs. Following Davidson’s model, F3-F4 receives a 1.2 attention weight.
Stream outputs are fused using learned softmax weights, with a learnable skip connection preserving convolutional features. The skip weight is initialized such that approximately 73% of the signal comes from the convolutional path. We explored three model sizes (138K, 419K, and 1.6M parameters), with the smallest recommended for this dataset size.
4.2.2 M2-EEG: EEGNet-Improved
As a literature-based alternative, we enhanced EEGNet with techniques from recent papers: squeeze-excitation blocks for channel attention, multi-scale temporal convolution (kernels 16, 32, 64), residual connections, data augmentation (amplitude scaling, Gaussian noise, time shift), label smoothing (0.1), AdamW optimizer with weight decay, and cosine annealing learning rate schedule.
4.2.3 M2-Audio: Temporal-Frequency Dual Attention
Audio spectrograms have natural 2D structure: the temporal axis captures speech rhythm, pauses, and intensity changes, while the frequency axis captures pitch, harmonics, and voice quality. Standard AST flattens this structure into a 1D patch sequence, losing the distinction.
Our architecture preserves AST’s pretrained features but adds factorized attention. After extracting 1212 patches from AST, we reshape them into a 10112 grid (timefrequency). Learnable 2D positional encodings are added. Two dual-attention blocks apply temporal attention (each frequency band attends across time) followed by frequency attention (each time step attends across frequencies). A learnable skip connection, initialized with weight (yielding ), blends the attention output with AST’s original CLS token, preserving approximately 88% of pretrained features.
4.2.4 M2-Vision: Factorized Space-Time Attention
Video understanding requires both spatial analysis within frames and temporal modeling across frames. Full space-time attention over 25 frames 196 patches has prohibitive million attention entries.
Following ViViT’s Model 2 (factorized encoder), we decompose this into spatial attention within each frame () followed by temporal attention for each patch position across frames (), achieving 22 speedup. Four factorized attention blocks process ViT’s patch embeddings (shape: batch25196768), with separate learnable spatial and temporal positional encodings. A skip connection preserves ViT’s pretrained features.
4.3 M3: CNN Improvements
After M2’s poor results, we investigated whether simple fixes to baseline CNNs could outperform complex attention mechanisms.
4.3.1 M3-Audio: Delta MFCCs
The original audio CNN uses 180 features (40 MFCCs, 12 chroma, 128 mel), averaged over time. We hypothesized that adding temporal dynamics without dramatically increasing dimensionality would help. Delta MFCCs capture the first derivative of cepstral coefficients, encoding how spectral features change over time, which is important for emotional speech where pitch and intensity vary dynamically.
Our modification adds 40 delta MFCC features (also time-averaged), increasing the feature vector from 180 to 220 dimensions. Combined with AdamW optimizer (weight decay 1e-4) and label smoothing (0.1), this minimal change achieved our best audio result.
4.3.2 M3-EEG v1: Bug Fixes
We identified a bug in the original EEGNet implementation: nn.Softmax(dim=1) in the classifier output, combined with CrossEntropyLoss which applies log-softmax internally. This computes , damaging gradients. We removed the redundant softmax, added AdamW optimizer with weight decay (0.01), and label smoothing (0.1).
4.3.3 M3-EEG v2: Band Power Features
Bug fixes alone proved insufficient because raw time-domain EEG has poor signal-to-noise ratio: approximately 80% of the signal consists of artifacts from muscle activity, eye blinks, and electrical noise. Following established practices from SEED dataset research Zheng and Lu (2015), we extracted frequency-domain features.
Band power features compute power spectral density using Welch’s method for five frequency bands (delta: 0.5 to 4 Hz, theta: 4 to 8 Hz, alpha: 8 to 13 Hz, beta: 13 to 30 Hz, gamma: 30 to 45 Hz) across all 30 channels, yielding 150 features. Differential entropy features compute for each band-filtered signal, adding another 150 features. Alpha asymmetry features compute for six hemisphere pairs, adding 6 features that capture frontal asymmetry.
The 306-dimensional feature vector feeds a simple MLP (306128645) with batch normalization and 50% dropout. Despite having more parameters than EEGNet, this approach operates on meaningful, low-noise features rather than raw high-dimensional signals.
Concrete Example.
Raw EEG input has 15,000 values (30 channels 500 time points) with approximately 20% signal and 80% noise. Band power features compress this to 306 values by: (1) transforming to frequency domain, which averages out temporal noise; (2) focusing on emotion-relevant bands (alpha, beta); and (3) computing asymmetry features grounded in neuroscience. A neural network would need to learn this entire transformation from 280 samples, which is not feasible.
4.3.4 M3-Vision v1: SE Block Fix
We discovered that the original video CNN’s squeeze-excitation block used reduction ratio 1, meaning no dimensionality reduction occurred (204820482048 instead of 20481282048). This defeats the purpose of SE blocks, which learn channel importance through a bottleneck Hu et al. (2018). We fixed the ratio to 16.
4.3.5 M3-Vision v2: Delta Features
Following our audio success, we added temporal delta features to the video CNN, capturing frame-to-frame changes in facial expression. Delta features compute the difference between consecutive frame representations, encoding expression dynamics that static features miss.
5 Experiments and Results
All models were trained on NVIDIA RTX 6000 Ada GPUs (48GB VRAM). We report mean accuracy and standard deviation across all 42 subjects.
| Category | Model | EEG | Audio | Vision |
| Paper Baselines | CNN | 60.00% | 61.90% | 71.40% |
| Transformer | 53.50% | 62.70% | 74.50% | |
| M1: Our Baseline | Transformer | 52.68% 10.3 | 58.06% 10.5 | 75.30% 10.5 |
| M2: Factorized Attn. | EEG Tri-Stream | 48.00% 12.1 | — | — |
| EEGNet-Improved | 58.47% 11.2 | — | — | |
| Audio Dual-Attn. | — | 49.46% 15.8 | — | |
| Vision ViViT-style | — | — | 69.54% 12.3 | |
| M3: CNN Improved | EEG v1 (Bug Fix) | 57.66% 13.7 | — | — |
| EEG v2 (Band Power) | 67.62% 8.2 | — | — | |
| Audio (Delta MFCC) | — | 65.56% 9.3 | — | |
| Vision v1 (SE Fix) | — | — | 69.42% 11.9 | |
| Vision v2 (Delta) | — | — | 72.68% 10.8 |
| Subject | EEG | Audio | Vision |
| Top Performers | |||
| S17 | 82.50% | 85.00% | 93.33% |
| S6 | — | 78.33% | 94.17% |
| S20 | 83.33% | — | 91.67% |
| S24 | 85.00% | — | — |
| S2 | 80.83% | 85.00% | — |
| Difficult Subjects | |||
| S35 | 46.67% | 55.00% | 56.67% |
| S12 | — | 54.17% | 43.33% |
| S29 | 56.67% | 47.50% | — |
| S37 | — | 46.67% | — |
| S34 | 54.17% | 49.17% | — |
Table 1 presents comprehensive results. Three findings emerge clearly.
M2 Factorized Attention Fails.
All M2 models underperform their respective baselines. The EEG Tri-Stream achieves only 48% (vs. 60% paper baseline), Audio Dual-Attention reaches 49.46% with 7 subjects stuck at random chance (20%), and Vision ViViT-style achieves 69.54% (vs. 74.5% paper baseline). Despite theoretically sound motivation, these complex architectures cannot learn effectively from 280 training samples.
M3 Simple Fixes Succeed.
Domain-appropriate features improve performance substantially. EEG Band Power achieves 67.62% (+7.62pp over paper), Audio Delta MFCCs achieves 65.56% (+3.66pp), and Vision Delta features achieve 72.68% (+1.28pp over paper CNN). Established signal processing techniques (frequency analysis, differential features) extract more emotion-relevant information than complex attention mechanisms can learn from limited data. Figure 2 illustrates these architectures.
Vision: Pretraining Domain Matters.
For Vision, M1’s domain-specific pretraining (75.30%) outperformed all M3 CNN modifications (72.68%), demonstrating that pretraining domain matters more than architectural changes. The ViT model pretrained on facial emotions provides better features than ResNet50 pretrained on ImageNet, even with our improvements.
Subject Variability.
Table 2 shows substantial individual differences. Subject 17 consistently achieves highest accuracy across all modalities (93.33% vision, 85% audio, 82.5% EEG), suggesting highly expressive emotional displays. Subjects 35 and 12 present persistent challenges, potentially due to subtle expressions or individual physiological differences.
6 Analysis: Why M2 Failed
Our M2 architectures failed for three interconnected reasons.
Destruction of Pretrained Features.
AST and ViT backbones contain valuable features learned from millions of images and audio clips. Our additional attention layers, despite skip connections, modify these representations. The learned skip weight () indicates the model tries to preserve pretrained features (88%), but this proves insufficient: the 12% attention contribution still degrades performance.
Insufficient Training Data.
The EEG Tri-Stream model has 138,401 parameters trained on 280 samples per subject. Standard heuristics suggest parameters should be less than 10 training samples for good generalization. Our models exceed this by 5 to 50, leading to overfitting (100% training accuracy but 20 to 80% test accuracy, especially for Audio M2 where 7 subjects achieve exactly 20% random chance).
Complex Inductive Biases Require Data.
Factorized attention implements sophisticated inductive biases (separate spatial, temporal, and frequency reasoning), but learning to use these biases effectively requires substantial data. On small datasets, the overhead of additional parameters outweighs any architectural benefit.
7 Analysis: Why M3 Succeeded
Domain Knowledge Over Learned Features.
Band power features encode decades of neuroscience research into 306 meaningful dimensions. Alpha asymmetry directly captures Davidson’s frontal asymmetry model. Delta MFCCs encode speech dynamics known to correlate with emotion. These features compress relevant information while discarding noise, something neural networks struggle to learn from limited data.
Bug Fixes Restore Intended Behavior.
The softmax bug in EEGNet and SE ratio bug in the video CNN were non-obvious implementation errors that degraded baseline performance. Simple fixes recovered significant accuracy without adding complexity.
Minimal Complexity.
M3 models add minimal parameters over baselines. The 220-feature audio CNN adds only 40 dimensions. The band power MLP, while larger than EEGNet, operates on preprocessed features rather than raw 15,000-dimensional signals.
8 Limitations
Our study has several limitations. First, all experiments use subject-dependent evaluation; cross-subject generalization remains unexplored. Second, the EAV dataset’s conversational paradigm may not generalize to spontaneous emotions. Third, we did not explore multimodal fusion, which may further improve results by combining complementary information across modalities. Fourth, our M2 architectures may benefit from more careful hyperparameter tuning on larger validation sets. Finally, the EAV dataset contains only 42 subjects, limiting statistical power for detecting small effects.
9 Conclusion
This study systematically investigated attention mechanisms for multimodal emotion recognition on small datasets. Our findings show that for datasets with approximately 280 training samples per subject, domain-appropriate signal processing outperforms sophisticated neural architectures.
Our best results (EEG 67.62%, +7.62pp; Audio 65.56%, +3.66pp; Vision 75.30%, +0.80pp transformer and 72.68%, +1.28pp CNN) were achieved through simple modifications: frequency-domain features for EEG, delta coefficients for audio and video, and properly pretrained transformers. Meanwhile, our theoretically motivated factorized attention mechanisms (M2) consistently underperformed by 5 to 13 percentage points.
The practical lesson is clear: before adding architectural complexity, researchers should (1) verify baseline implementations for bugs, (2) consider domain-specific feature engineering, and (3) match model capacity to available training data. Future work should evaluate these factorized attention approaches on larger emotion datasets where their sophisticated inductive biases may prove beneficial.
Appendix A Theoretical Background
This appendix provides comprehensive theoretical foundations for the deep learning models and signal processing techniques used in multimodal emotion recognition, organized by model category (M1, M2, M3).
A.1 A. Foundations
A.1.1 A.1 Emotion Recognition Fundamentals
Discrete Emotion Theory proposes that emotions consist of universal basic categories. The EAV dataset uses five discrete classes: Neutral, Anger, Happiness, Sadness, and Calmness. These categories map to Russell’s circumplex model of affect along valence (positive/negative) and arousal (high/low) dimensions:
| Happiness | (1) | |||
| Anger | ||||
| Sadness | ||||
| Calmness | ||||
| Neutral |
A.1.2 A.2 EEG Signal Processing
Electroencephalography (EEG) measures electrical activity from neural populations via scalp electrodes. The International 10-20 System specifies electrode placement, where letters indicate brain regions (F=Frontal, C=Central, P=Parietal, O=Occipital, T=Temporal) and numbers indicate hemisphere (odd=left, even=right, z=midline).
The EAV dataset uses 30 channels sampled at 500 Hz. Key electrode pairs for emotion include:
F3-F4: Frontal asymmetry (primary emotion indicator)
Fp1-Fp2: Prefrontal activity
C3-C4: Central sensorimotor
EEG signals are characterized by frequency bands with distinct functional correlates:
| Band | Frequency | Function |
| Delta () | 0.5–4 Hz | Deep sleep, unconscious processing |
| Theta () | 4–8 Hz | Memory encoding, emotional processing |
| Alpha () | 8–13 Hz | Relaxed alertness, frontal asymmetry |
| Beta () | 13–30 Hz | Active thinking, concentration |
| Gamma () | 30–45 Hz | High-level cognitive binding |
Davidson’s Frontal Asymmetry Model (1992) establishes that left frontal cortex activation (F3) correlates with approach motivation and positive emotions, while right frontal activation (F4) correlates with withdrawal motivation and negative emotions. The Frontal Asymmetry Index (FAI) quantifies this:
where denotes alpha band power. Note that alpha power is inversely related to cortical activity (higher alpha = lower activity), so positive FAI indicates greater left frontal activation (positive valence).
A.1.3 A.2.1 EEG Preprocessing Pipeline
Step 1: Downsampling. Original 500 Hz 100 Hz using polyphase resampling:
where is the decimation factor and is an anti-aliasing FIR filter.
Step 2: Bandpass Filtering. 5th-order Butterworth IIR filter (0.5–45 Hz):
where is the filter order. Transfer function:
A.1.4 A.3 Audio Signal Processing
Mel-Frequency Cepstral Coefficients (MFCCs) are the standard features for speech analysis, designed to mimic human auditory perception. The extraction pipeline consists of five steps:
Step 1: Pre-emphasis. Boost high frequencies to balance the spectrum:
Step 2: Framing and Windowing. Segment into overlapping frames with Hamming window:
Step 3: Short-Time Fourier Transform.
Step 4: Mel Filterbank. Apply triangular filters on mel scale:
Step 5: Discrete Cosine Transform. Decorrelate filterbank outputs:
Delta Features capture temporal dynamics:
A.1.5 A.4 Vision Processing
Facial Action Coding System (FACS) (Ekman & Friesen, 1978) decomposes facial expressions into Action Units (AUs)—minimal muscle movements:
| Emotion | Typical AUs | Description |
| Happiness | AU6 + AU12 | Cheek raiser + Lip corner puller (Duchenne smile) |
| Sadness | AU1 + AU4 + AU15 | Inner brow raiser + Brow lowerer + Lip corner depressor |
| Anger | AU4 + AU5 + AU7 | Brow lowerer + Upper lid raiser + Lid tightener |
| Fear | AU1 + AU2 + AU4 + AU20 | Brows raised, pulled together, lip stretch |
| Surprise | AU1 + AU2 + AU5 + AU26 | Brow raiser + Upper lid raiser + Jaw drop |
A.1.6 A.5 Transformer Architecture
Scaled Dot-Product Attention:
where:
scaling prevents softmax saturation
Multi-Head Attention:
Transformer Encoder Layer:
Computational complexity is , quadratic in sequence length .
A.1.7 A.6 CNN Architectures
EEGNet Architecture. Uses depthwise separable convolutions for efficient EEG processing:
Block 1: Temporal Convolution
where at 100Hz captures 3-second temporal patterns.
Depthwise Convolution: Learns spatial filters per channel
Block 2: Separable Convolution
Parameter Count: Only 2,600 parameters—ideal for small datasets!
Squeeze-and-Excitation Blocks. Learn channel-wise attention:
Squeeze: Global average pooling
Excitation: Bottleneck MLP with reduction ratio
where ,
Scale: Channel reweighting
Critical Bug in Original Code: means —no bottleneck, no squeeze! We fixed to .
A.1.8 A.7 Vision Transformer for Video
Patch Embedding:
For a image with patches:
Classification:
where is the [CLS] token output from the final layer.
A.2 B. Model 1: Baseline Transformers
M1 establishes baseline performance using standard transformer architectures with minimal modifications from pretrained models.
A.2.1 B.1 M1-EEG: Custom EEG Transformer
Our EEG baseline uses a lightweight custom architecture since no suitable pretrained EEG transformer exists for emotion recognition.
Convolutional Feature Extraction. Two 1D convolutions extract local temporal features:
Transformer Encoding. Six encoder layers with 4 attention heads process temporal dependencies:
Classification. Mean pooling followed by MLP:
Limitation: Standard self-attention treats all time points as a flat sequence, failing to separate spatial (channel) and temporal (time) patterns.
Result: 52.68% 10.28
A.2.2 B.2 M1-Audio: Audio Spectrogram Transformer (AST)
We fine-tune the pretrained AST model (MIT/ast-finetuned-audioset), which treats spectrograms as images using Vision Transformer architecture.
Patch Embedding. The spectrogram (1024128) is divided into 1616 patches:
Two-Stage Training.
1.
Stage 1: Freeze backbone, train classifier (LR=5e-4, 10 epochs)
2.
Stage 2: Unfreeze all, fine-tune (LR=5e-6, 15 epochs)
Limitation: Flattening patches loses the distinction between temporal (rhythm) and frequency (pitch) axes.
Result: 58.06% 10.54
A.2.3 B.3 M1-Vision: Vision Transformer (ViT)
We use a ViT pretrained on facial emotion recognition (dima806/facial_emotions_image_detection).
Frame-Level Processing. Each of 25 video frames is processed independently:
Video-Level Prediction. Frame predictions are averaged:
Limitation: No temporal modeling across frames—cannot capture facial dynamics.
Result: 75.30% 10.45 (exceeds paper baseline by +0.80pp)
A.3 C. Model 2: Factorized Attention
M2 represents our main architectural contribution: axis-specific factorized attention tailored to each modality’s unique structure.
A.3.1 C.1 Motivation: Why Factorized Attention?
Standard transformers flatten multi-dimensional data into 1D sequences, losing structural information. For emotion recognition:
EEG: Spatial (brain regions) vs temporal (dynamics) vs asymmetry (hemisphere differences)
Audio: Temporal (rhythm) vs frequency (pitch)
Video: Spatial (face regions) vs temporal (expression dynamics)
Factorized attention models these axes separately, providing better inductive bias and computational efficiency.
Complexity Analysis. For video with frames and patches:
For , :
A.3.2 C.2 M2-EEG v1: Tri-Stream Spatial-Temporal-Asymmetry Transformer
Novel Contribution: We introduce a third attention stream for hemispheric asymmetry, based on Davidson’s frontal asymmetry model.
Architecture. Input EEG (30 channels, 500 samples) passes through convolutional feature extraction, then branches into three parallel attention streams:
Spatial Stream. Channels attend to channels at each time point:
This captures which brain regions interact (functional connectivity).
Temporal Stream. Time points attend to time points for each channel:
This captures event-related dynamics.
Asymmetry Stream. We compute difference signals for six hemisphere pairs:
The F3-F4 pair receives a subtle 1.2 attention weight boost based on its established importance for emotional valence.
Fusion. Streams are combined via learned softmax weights:
Skip Connection. A learnable skip preserves convolutional features:
where with initialized to yield .
Model Variants:
| Size | Params | Layers | Heads |
| Small (recommended) | 138,401 | 2 | 2 |
| Base | 419,201 | 3 | 4 |
| Large | 1,626,305 | 4 | 8 |
Result: 48% (fails—too complex for 280 samples)
Failure Analysis: With 138K parameters and 280 training samples, the model has 50 more parameters than data points, leading to severe overfitting. The sophisticated tri-stream architecture cannot learn meaningful patterns from such limited data.
A.3.3 C.3 M2-EEG v2: EEGNet-Improved
As a literature-based alternative, we enhanced EEGNet with proven techniques:
Squeeze-Excitation Blocks. Learn channel importance:
Multi-Scale Temporal Convolution. Parallel convolutions with kernels [16, 32, 64] capture different temporal scales.
Residual Connections. Help gradient flow in deeper networks.
Additional Improvements: Label smoothing (0.1), AdamW with weight decay (1e-3), cosine annealing LR, data augmentation (amplitude scaling, Gaussian noise, time shift).
Result: 58.47% 12.56 (approaches paper baseline)
A.3.4 C.4 M2-Audio: Temporal-Frequency Dual Attention
Core Idea. Preserve AST’s pretrained features while adding axis-specific attention.
Grid Reconstruction. Reshape AST’s 1212 patch embeddings into a 10112 grid (time frequency):
2D Positional Encoding. Add learnable temporal and frequency position embeddings:
Dual Attention Blocks. Two blocks apply sequential attention:
Temporal Attention: Each frequency band attends across time:
Frequency Attention: Each time step attends across frequency:
Skip Connection. Critical for preserving AST features:
where initially (88% AST, 12% attention).
Result: 49.46% (fails—7 subjects stuck at 20% random chance)
Failure Analysis: Despite preserving 88% of AST features, the additional attention layers introduce noise that corrupts the pretrained representations. On small data, even minor corruption cannot be recovered.
A.3.5 C.5 M2-Vision: Factorized Space-Time Attention
Architecture. Following ViViT Model 2 (factorized encoder), we apply:
Spatial Attention. Within each frame, patches attend to patches:
Temporal Attention. For each patch position, frames attend to frames:
Skip Connection. Preserves ViT features with learnable .
Result: 69.54% (best M2 result, but still -5pp below baseline)
Partial Success Analysis: Skip connection helps preserve pretrained features. However, the additional spatial-temporal attention layers still degrade performance compared to simple frame averaging, suggesting that temporal modeling is unnecessary for this dataset’s short video clips.
A.4 D. Model 3: CNN Improvements
After M2’s failures, we investigated whether simple fixes to baseline CNNs could outperform complex attention mechanisms.
A.4.1 D.1 Philosophy: Simple Fixes Over Complex Architectures
M2’s failures revealed that:
1.
Complex models overfit on small datasets
2.
Pretrained features are destroyed by additional layers
3.
The baseline authors likely tried complex approaches first
M3 philosophy: Fix bugs, add minimal domain-appropriate features, preserve working architecture.
A.4.2 D.2 M3-Audio: Delta MFCCs
Bug Fix: Replace Adam with AdamW (proper weight decay).
Minimal Addition: Add 40 delta MFCC features (180 220 total):
Why Not Temporal Features? Keeping full temporal MFCCs (40100 = 4000 values) would cause massive overfitting with 280 samples. Time-averaged delta features capture dynamics without dimensional explosion.
Training: AdamW (weight decay 1e-4), label smoothing (0.1).
Result: 65.56% 9.30 (+3.66pp over paper baseline)
A.4.3 D.3 M3-EEG v1: Bug Fixes Only
Critical Bug Found: The original EEGNet classifier contained nn.Softmax(dim=1) before the loss function. Combined with CrossEntropyLoss (which applies log-softmax internally), this computes:
This double softmax severely damages gradients—the outer softmax on already-normalized probabilities produces near-uniform distributions, destroying learning signal.
Fix: Remove nn.Softmax from classifier, use raw logits with CrossEntropyLoss.
Additional Changes: AdamW (weight decay 0.01), label smoothing (0.1).
Result: 57.66% 13.66 (-2.34pp below paper baseline)
Analysis: Bug fix alone insufficient. Raw time-domain EEG has poor signal-to-noise ratio (80% artifacts from muscle, eye, and electrical noise).
A.4.4 D.4 M3-EEG v2: Band Power Features (Best EEG Result)
Key Insight: Instead of learning frequency decomposition from raw EEG (15,000 values), extract established frequency-domain features first.
Feature Extraction Pipeline:
1. Band Power (150 features). For each channel and frequency band , compute power spectral density using Welch’s method:
where is estimated via:
Features: 30 channels 5 bands = 150.
2. Differential Entropy (150 features). For each band-filtered signal assumed Gaussian:
This is standard in SEED dataset literature. Features: 30 5 = 150.
3. Alpha Asymmetry (6 features). Following Davidson’s model:
for pairs (F3,F4), (Fp1,Fp2), (F7,F8), (C3,C4), (P3,P4), (O1,O2).
Total: 306 features.
Classifier: Simple MLP:
Why This Works:
| Aspect | Raw EEG (v1) | Band Power (v2) |
| Input dimension | 15,000 | 306 |
| Signal-to-noise | 20% | High (averaged) |
| Frequency info | Must learn | Pre-extracted |
| Asymmetry info | Not explicit | 6 dedicated features |
Result: 67.62% 7.89 (+7.62pp over paper baseline)
A.4.5 D.5 M3-Vision v1: SE Block Fix
Bug Found: The original squeeze-excitation block used reduction ratio :
With , there is no bottleneck—the “squeeze” does not squeeze! The SE block degenerates to a pointless transformation.
Fix: Set for proper bottleneck:
For : (proper squeeze-excitation).
Result: 69.42% 10.81 (-1.98pp vs paper baseline)
A.4.6 D.6 M3-Vision v2: Delta Features (Best CNN)
Following audio success, we add frame-to-frame delta features capturing expression dynamics. The delta features compute the difference between consecutive frame representations:
where is the frame feature vector from the backbone.
The final representation combines averaged static and delta features:
Result: 72.68% 9.56 (+1.28pp over paper CNN baseline)
A.5 E. Optimization and Regularization
A.5.1 E.1 Loss Functions
Cross-Entropy Loss:
Label Smoothing. Softens targets to prevent overconfidence:
With and :
A.5.2 E.2 AdamW Optimizer
Key difference from Adam: weight decay is applied directly, not scaled by adaptive learning rate.
A.5.3 E.3 Regularization
Dropout: Randomly zeros elements with probability :
Batch Normalization:
Gradient Clipping: Prevents exploding gradients:
A.5.4 E.4 Skip Connections in M2
Critical for preserving pretrained features:
Initialization: with yields (88% pretrained, 12% new).
Observation: After training, remains near 0.12—the model learns that pretrained features dominate on small data.
A.6 F. Model Complexity Analysis
| Model | Params | Trainable | Result |
| EEG Models | |||
| EEGNet (Paper) | 2,600 | 2,600 | 60.0% |
| M3-EEG v1 (Bug Fix) | 2,600 | 2,600 | 57.66% |
| M3-EEG v2 (Band Power) | 48K | 48K | 67.62% |
| M2-EEGNet-Improved | 36,501 | 36,501 | 58.47% |
| M2-Tri-Stream (small) | 138,401 | 138,401 | 48% |
| Audio Models | |||
| AST (M1) | 87M | 4K | 58.06% |
| M3-Audio (Delta) | 87M | 5K | 65.56% |
| M2-Dual Attention | 87M | 12M | 49.46% |
| Vision Models | |||
| ViT (M1) | 86M | 4K | 75.30% |
| M3-Vision v1 (SE Fix) | 23M | 23M | 69.42% |
| M3-Vision v2 (Delta) | 23M | 23M | 72.68% |
| M2-Factorized | 86M | 12M | 69.54% |
Rule of Thumb: Trainable parameters should be 10 training samples. With 280 samples, this suggests 2,800 parameters for optimal generalization.
Observation: M3-EEG v2 succeeds despite 48K parameters because it operates on meaningful 306-dimensional features rather than noisy 15,000-dimensional raw signals.
A.7 G. Data Augmentation
EEG:
Amplitude scaling: ,
Gaussian noise: ,
Time shift: Circular shift by 10%
Audio (SpecAugment):
Time masking: Zero consecutive frames,
Frequency masking: Zero consecutive bins,
Video:
Horizontal flip with
Color jitter (brightness, contrast, saturation)
Random erasing (simulates occlusion)
A.8 H. Evaluation Metrics
A.8.1 H.1 Accuracy
A.8.2 H.2 F1 Score (Weighted)
Per-class metrics:
F1 Score:
Weighted Average:
where is the number of samples in class .
A.9 I. Analysis: Why Methods Failed or Succeeded
A.9.1 I.1 Why M2 Factorized Attention Failed
1.
Destruction of Pretrained Features: AST and ViT backbones contain valuable features learned from millions of samples. Our attention layers, despite skip connections (), still corrupt these representations.
2.
Insufficient Training Data: With 280 samples and 138K+ parameters, M2 models have 50 more parameters than data points. Standard heuristics suggest 10 for good generalization.
3.
Complex Inductive Biases Require Data: Factorized attention implements sophisticated biases (separate spatial/temporal reasoning), but learning to use them requires substantial data.
Evidence: M2 Audio had 7/42 subjects stuck at exactly 20% (random chance for 5 classes), indicating complete failure to learn.
A.9.2 I.2 Why M3 Simple Fixes Succeeded
1.
Domain Knowledge Over Learned Features: Band power features encode decades of neuroscience research into 306 meaningful dimensions. Delta MFCCs encode known speech dynamics.
2.
Bug Fixes Restore Intended Behavior: The softmax bug and SE ratio bug were non-obvious errors degrading baseline performance.
3.
Minimal Complexity: M3 adds minimal parameters. The 220-feature audio CNN adds only 40 dimensions over baseline.
Key Insight: With 280 training samples, fixing bugs is more valuable than adding features.

### v2

@@
 The EAV dataset Kim et al.
-(2024) provides synchronized EEG, Audio, and Video recordings from 42 participants engaged in conversational interactions, offering a unique resource for multimodal emotion research.
+(2024) provides synchronized EEG, Audio, and Video recordings from 42 participants engaged in conversational interactions, offering a unique resource for multimodal emotion research alongside large-scale multimodal affect datasets such as CMU-MOSEI Bagher Zadeh et al.
+(2018).
 However, with only approximately 280 training samples per subject after data splitting, this dataset presents significant challenges for deep learning approaches.
@@
 Traditional EEG datasets such as DEAP Koelstra et al.
-(2012) and MAHNOB-HCI employ passive elicitation paradigms where participants watch emotional stimuli.
+(2012) and MAHNOB-HCI Soleymani et al.
+(2012) employ passive elicitation paradigms where participants watch emotional stimuli.
 Audio-visual datasets like IEMOCAP Busso et al.
@@
 Classical approaches to emotion recognition rely on hand-crafted features.
-For EEG, band power and differential entropy features computed from frequency bands (delta, theta, alpha, beta, gamma) have proven effective on the SEED dataset Zheng and Lu (2015).
-For audio, Mel-frequency cepstral coefficients (MFCCs) and their temporal derivatives (delta and delta-delta) capture both spectral content and dynamics Schuller and others (2013).
+For EEG, band power and differential entropy features computed from frequency bands (delta, theta, alpha, beta, gamma) have proven effective on the SEED dataset Zheng and Lu (2015) and remain widely used across EEG emotion recognition studies Patel et al.
+(2021).
+For audio, Mel-frequency cepstral coefficients (MFCCs) and their temporal derivatives (delta and delta-delta) capture both spectral content and dynamics Schuller and others (2013); El Ayadi et al.
+(2011); Ververidis and Kotropoulos (2006), and are commonly extracted using toolkits such as openSMILE Eyben et al.
+(2010).
 Squeeze-and-Excitation networks Hu et al.
@@
 The same two-stage training strategy is applied: 10 epochs frozen, then 5 epochs unfrozen.
-While this approach ignores temporal dynamics across frames, the facial emotion pretraining provides strong baseline performance.
+While this approach ignores temporal dynamics across frames, the facial emotion pretraining (often built from large-scale facial affect data such as AffectNet) provides strong baseline performance Mollahosseini et al.
+(2017).
 4.2 M2: Factorized Attention Mechanisms
@@
 The original audio CNN uses 180 features (40 MFCCs, 12 chroma, 128 mel), averaged over time.
-We hypothesized that adding temporal dynamics without dramatically increasing dimensionality would help.
-Delta MFCCs capture the first derivative of cepstral coefficients, encoding how spectral features change over time, which is important for emotional speech where pitch and intensity vary dynamically.
+To introduce temporal dynamics with minimal additional capacity, we augment the feature set with delta MFCCs.
+Delta MFCCs approximate the first temporal derivative of the cepstral coefficients and capture short-term spectral change patterns that are informative for affective speech (e.g., variations in pitch and energy).
 Our modification adds 40 delta MFCC features (also time-averaged), increasing the feature vector from 180 to 220 dimensions.
@@
 4.3.2 M3-EEG v1: Bug Fixes
-We identified a bug in the original EEGNet implementation: nn.Softmax(dim=1) in the classifier output, combined with CrossEntropyLoss which applies log-softmax internally.
-This computes , damaging gradients.
-We removed the redundant softmax, added AdamW optimizer with weight decay (0.01), and label smoothing (0.1).
+We identified an implementation issue in the original EEGNet classifier: applying nn.Softmax(dim=1) before CrossEntropyLoss, which already applies (log-)softmax internally.
+This double normalization can distort gradients and hinder optimization.
+We therefore removed the explicit softmax and trained the model on raw logits, using AdamW with weight decay (0.01) and label smoothing (0.1).
 4.3.3 M3-EEG v2: Band Power Features
-Bug fixes alone proved insufficient because raw time-domain EEG has poor signal-to-noise ratio: approximately 80% of the signal consists of artifacts from muscle activity, eye blinks, and electrical noise.
-Following established practices from SEED dataset research Zheng and Lu (2015), we extracted frequency-domain features.
+Bug fixes alone proved insufficient, as raw time-domain EEG is often heavily contaminated by artifacts (e.g., muscle activity, eye blinks, and environmental noise).
+Following established practice in EEG emotion recognition (e.g., SEED-style pipelines) Zheng and Lu (2015), we therefore extracted frequency-domain features.
 Band power features compute power spectral density using Welch’s method for five frequency bands (delta: 0.5 to 4 Hz, theta: 4 to 8 Hz, alpha: 8 to 13 Hz, beta: 13 to 30 Hz, gamma: 30 to 45 Hz) across all 30 channels, yielding 150 features.
@@
 Despite having more parameters than EEGNet, this approach operates on meaningful, low-noise features rather than raw high-dimensional signals.
-Concrete Example.
-Raw EEG input has 15,000 values (30 channels 500 time points) with approximately 20% signal and 80% noise.
-Band power features compress this to 306 values by: (1) transforming to frequency domain, which averages out temporal noise; (2) focusing on emotion-relevant bands (alpha, beta); and (3) computing asymmetry features grounded in neuroscience.
-A neural network would need to learn this entire transformation from 280 samples, which is not feasible.
+Illustrative Comparison.
+The raw EEG segment contains 15,000 samples (30 channels 500 time points) and includes substantial nuisance variability.
+In contrast, the 306-dimensional feature representation applies (i) a frequency-domain transformation with averaging (Welch), (ii) band selection that emphasizes physiologically meaningful rhythms (e.g., alpha/beta), and (iii) explicit hemispheric asymmetry features.
+Learning an equivalent denoising and summarization pipeline end-to-end from 280 training samples per subject is unlikely to generalize.
 4.3.4 M3-Vision v1: SE Block Fix
-We discovered that the original video CNN’s squeeze-excitation block used reduction ratio 1, meaning no dimensionality reduction occurred (204820482048 instead of 20481282048).
-This defeats the purpose of SE blocks, which learn channel importance through a bottleneck Hu et al.
-(2018).
-We fixed the ratio to 16.
+We found that the squeeze-excitation (SE) module in the original video CNN used reduction ratio 1, i.e., no bottleneck (204820482048 rather than 20481282048).
+Since SE relies on a bottleneck to learn channel reweighting efficiently Hu et al.
+(2018), we set the reduction ratio to 16.
 4.3.5 M3-Vision v2: Delta Features
@@
 5 Experiments and Results
-All models were trained on NVIDIA RTX 6000 Ada GPUs (48GB VRAM). We report mean accuracy and standard deviation across all 42 subjects.
+All models were trained on NVIDIA RTX 6000 Ada GPUs with 48GB of VRAM. We report mean accuracy and standard deviation across all 42 subjects.
 | Category | Model | EEG | Audio | Vision |
@@
 | S34 | 54.17% | 49.17% | — |
-Table 1 presents comprehensive results. Three findings emerge clearly.
-M2 Factorized Attention Fails.
-All M2 models underperform their respective baselines.
-The EEG Tri-Stream achieves only 48% (vs.
-60% paper baseline), Audio Dual-Attention reaches 49.46% with 7 subjects stuck at random chance (20%), and Vision ViViT-style achieves 69.54% (vs.
-74.5% paper baseline).
-Despite theoretically sound motivation, these complex architectures cannot learn effectively from 280 training samples.
-M3 Simple Fixes Succeed.
-Domain-appropriate features improve performance substantially.
-EEG Band Power achieves 67.62% (+7.62pp over paper), Audio Delta MFCCs achieves 65.56% (+3.66pp), and Vision Delta features achieve 72.68% (+1.28pp over paper CNN).
-Established signal processing techniques (frequency analysis, differential features) extract more emotion-relevant information than complex attention mechanisms can learn from limited data.
-Figure 2 illustrates these architectures.
+Table 1 summarizes results across all models. Three findings emerge.
+M2 Factorized Attention Underperforms.
+All M2 models fall below their respective baselines.
+The EEG Tri-Stream reaches 48% (vs.
+a 60% paper baseline), Audio Dual-Attention reaches 49.46% (with several subjects near chance-level performance), and the ViViT-style video model reaches 69.54% (vs.
+74.5%).
+Although the factorized designs are well motivated, they appear difficult to fit reliably in the low-sample regime ( 280 training samples per subject).
+M3 Simple Fixes Improve Performance.
+Domain-appropriate feature design yields consistent gains.
+EEG band-power features reach 67.62% (+7.62pp over the paper baseline), Audio delta MFCCs reach 65.56% (+3.66pp), and Vision delta features reach 72.68% (+1.28pp over the paper CNN).
+These results suggest that, under limited supervision, established signal-processing priors (frequency analysis and temporal derivatives) can provide more reliable inductive bias than additional attention modules.
+Figure 2 summarizes these models.
 Vision: Pretraining Domain Matters.
@@
 Table 2 shows substantial individual differences.
-Subject 17 consistently achieves highest accuracy across all modalities (93.33% vision, 85% audio, 82.5% EEG), suggesting highly expressive emotional displays.
+Subject 17 consistently achieves the highest accuracy across all modalities (93.33% vision, 85% audio, 82.5% EEG), suggesting relatively expressive emotional displays.
 Subjects 35 and 12 present persistent challenges, potentially due to subtle expressions or individual physiological differences.
-6 Analysis: Why M2 Failed
-Our M2 architectures failed for three interconnected reasons.
-Destruction of Pretrained Features.
-AST and ViT backbones contain valuable features learned from millions of images and audio clips.
-Our additional attention layers, despite skip connections, modify these representations.
-The learned skip weight () indicates the model tries to preserve pretrained features (88%), but this proves insufficient: the 12% attention contribution still degrades performance.
+6 Analysis: Why M2 Underperformed
+Our M2 architectures likely underperformed for three interconnected reasons.
+Degradation of Pretrained Representations.
+AST and ViT backbones encode strong representations learned from large-scale pretraining.
+Although we include skip connections, the additional attention layers still perturb these features.
+The learned skip weight () indicates a strong preference for the pretrained pathway ( 88%), yet the residual attention contribution can still negatively affect downstream accuracy.
 Insufficient Training Data.
-The EEG Tri-Stream model has 138,401 parameters trained on 280 samples per subject.
-Standard heuristics suggest parameters should be less than 10 training samples for good generalization.
-Our models exceed this by 5 to 50, leading to overfitting (100% training accuracy but 20 to 80% test accuracy, especially for Audio M2 where 7 subjects achieve exactly 20% random chance).
+The EEG Tri-Stream model has 138,401 parameters but only 280 training samples per subject.
+In this setting, model capacity can easily exceed what the data can constrain, leading to high variance and unstable generalization.
+Empirically, we observed strong overfitting (near-perfect training accuracy with substantially lower test accuracy), particularly for the M2 audio variant where multiple subjects collapsed toward chance-level performance.
 Complex Inductive Biases Require Data.
-Factorized attention implements sophisticated inductive biases (separate spatial, temporal, and frequency reasoning), but learning to use these biases effectively requires substantial data.
-On small datasets, the overhead of additional parameters outweighs any architectural benefit.
+Factorized attention introduces additional structure (e.g., separate spatial, temporal, and frequency reasoning).
+However, exploiting these inductive biases still requires sufficient supervision to estimate the extra parameters and to learn stable fusion.
+With limited data, the added flexibility can dominate any benefit and reduce robustness.
 7 Analysis: Why M3 Succeeded
 Domain Knowledge Over Learned Features.
-Band power features encode decades of neuroscience research into 306 meaningful dimensions.
-Alpha asymmetry directly captures Davidson’s frontal asymmetry model.
-Delta MFCCs encode speech dynamics known to correlate with emotion.
-These features compress relevant information while discarding noise, something neural networks struggle to learn from limited data.
-Bug Fixes Restore Intended Behavior.
-The softmax bug in EEGNet and SE ratio bug in the video CNN were non-obvious implementation errors that degraded baseline performance.
-Simple fixes recovered significant accuracy without adding complexity.
+Band power features incorporate well-established neurophysiological priors into a compact 306-dimensional representation.
+Alpha asymmetry directly operationalizes Davidson’s frontal asymmetry model.
+Delta MFCCs capture short-term speech dynamics that are commonly associated with affect.
+These features emphasize signal components that are difficult to learn reliably from limited data.
+Implementation Fidelity Dominates in the Low-Sample Regime.
+With only 280 training samples per subject, small implementation mismatches can materially change the effective learning problem and overwhelm the benefits of additional architectural components.
+We therefore performed targeted consistency checks that preserve the intended objective and inductive bias: (i) ensuring loss–activation compatibility by training on raw logits with CrossEntropyLoss (avoiding redundant softmax operations that distort gradients), and (ii) enforcing a true squeeze-and-excitation bottleneck via a non-trivial reduction ratio (so the module performs channel reweighting rather than a near-identity mapping).
+These corrections improve methodological validity and, in several cases, recover accuracy without increasing model capacity.
 Minimal Complexity.
@@
 Future work should evaluate these factorized attention approaches on larger emotion datasets where their sophisticated inductive biases may prove beneficial.
-Appendix A Theoretical Background
-This appendix provides comprehensive theoretical foundations for the deep learning models and signal processing techniques used in multimodal emotion recognition, organized by model category (M1, M2, M3).
-A.1 A. Foundations
-A.1.1 A.1 Emotion Recognition Fundamentals
-Discrete Emotion Theory proposes that emotions consist of universal basic categories.
-The EAV dataset uses five discrete classes: Neutral, Anger, Happiness, Sadness, and Calmness.
-These categories map to Russell’s circumplex model of affect along valence (positive/negative) and arousal (high/low) dimensions:
-| Happiness | (1) | |||
-| Anger | ||||
-| Sadness | ||||
-| Calmness | ||||
-| Neutral |
-A.1.2 A.2 EEG Signal Processing
-Electroencephalography (EEG) measures electrical activity from neural populations via scalp electrodes.
-The International 10-20 System specifies electrode placement, where letters indicate brain regions (F=Frontal, C=Central, P=Parietal, O=Occipital, T=Temporal) and numbers indicate hemisphere (odd=left, even=right, z=midline).
-The EAV dataset uses 30 channels sampled at 500 Hz. Key electrode pairs for emotion include:
-F3-F4: Frontal asymmetry (primary emotion indicator)
-Fp1-Fp2: Prefrontal activity
-C3-C4: Central sensorimotor
-EEG signals are characterized by frequency bands with distinct functional correlates:
-| Band | Frequency | Function |
-| Delta () | 0.5–4 Hz | Deep sleep, unconscious processing |
-| Theta () | 4–8 Hz | Memory encoding, emotional processing |
-| Alpha () | 8–13 Hz | Relaxed alertness, frontal asymmetry |
-| Beta () | 13–30 Hz | Active thinking, concentration |
-| Gamma () | 30–45 Hz | High-level cognitive binding |
-Davidson’s Frontal Asymmetry Model (1992) establishes that left frontal cortex activation (F3) correlates with approach motivation and positive emotions, while right frontal activation (F4) correlates with withdrawal motivation and negative emotions.
-The Frontal Asymmetry Index (FAI) quantifies this:
-where denotes alpha band power.
-Note that alpha power is inversely related to cortical activity (higher alpha = lower activity), so positive FAI indicates greater left frontal activation (positive valence).
-A.1.3 A.2.1 EEG Preprocessing Pipeline
-Step 1: Downsampling. Original 500 Hz 100 Hz using polyphase resampling:
-where is the decimation factor and is an anti-aliasing FIR filter.
-Step 2: Bandpass Filtering. 5th-order Butterworth IIR filter (0.5–45 Hz):
-where is the filter order. Transfer function:
-A.1.4 A.3 Audio Signal Processing
-Mel-Frequency Cepstral Coefficients (MFCCs) are the standard features for speech analysis, designed to mimic human auditory perception. The extraction pipeline consists of five steps:
-Step 1: Pre-emphasis. Boost high frequencies to balance the spectrum:
-Step 2: Framing and Windowing. Segment into overlapping frames with Hamming window:
-Step 3: Short-Time Fourier Transform.
-Step 4: Mel Filterbank. Apply triangular filters on mel scale:
-Step 5: Discrete Cosine Transform. Decorrelate filterbank outputs:
-Delta Features capture temporal dynamics:
-A.1.5 A.4 Vision Processing
-Facial Action Coding System (FACS) (Ekman & Friesen, 1978) decomposes facial expressions into Action Units (AUs)—minimal muscle movements:
-| Emotion | Typical AUs | Description |
-| Happiness | AU6 + AU12 | Cheek raiser + Lip corner puller (Duchenne smile) |
-| Sadness | AU1 + AU4 + AU15 | Inner brow raiser + Brow lowerer + Lip corner depressor |
-| Anger | AU4 + AU5 + AU7 | Brow lowerer + Upper lid raiser + Lid tightener |
-| Fear | AU1 + AU2 + AU4 + AU20 | Brows raised, pulled together, lip stretch |
-| Surprise | AU1 + AU2 + AU5 + AU26 | Brow raiser + Upper lid raiser + Jaw drop |
-A.1.6 A.5 Transformer Architecture
-Scaled Dot-Product Attention:
-where:
-scaling prevents softmax saturation
-Multi-Head Attention:
-Transformer Encoder Layer:
-Computational complexity is , quadratic in sequence length .
-A.1.7 A.6 CNN Architectures
-EEGNet Architecture. Uses depthwise separable convolutions for efficient EEG processing:
-Block 1: Temporal Convolution
-where at 100Hz captures 3-second temporal patterns.
-Depthwise Convolution: Learns spatial filters per channel
-Block 2: Separable Convolution
-Parameter Count: Only 2,600 parameters—ideal for small datasets!
-Squeeze-and-Excitation Blocks. Learn channel-wise attention:
-Squeeze: Global average pooling
-Excitation: Bottleneck MLP with reduction ratio
-where ,
-Scale: Channel reweighting
-Critical Bug in Original Code: means —no bottleneck, no squeeze! We fixed to .
-A.1.8 A.7 Vision Transformer for Video
-Patch Embedding:
-For a image with patches:
-Classification:
-where is the [CLS] token output from the final layer.
-A.2 B. Model 1: Baseline Transformers
-M1 establishes baseline performance using standard transformer architectures with minimal modifications from pretrained models.
-A.2.1 B.1 M1-EEG: Custom EEG Transformer
-Our EEG baseline uses a lightweight custom architecture since no suitable pretrained EEG transformer exists for emotion recognition.
-Convolutional Feature Extraction. Two 1D convolutions extract local temporal features:
-Transformer Encoding. Six encoder layers with 4 attention heads process temporal dependencies:
-Classification. Mean pooling followed by MLP:
-Limitation: Standard self-attention treats all time points as a flat sequence, failing to separate spatial (channel) and temporal (time) patterns.
-Result: 52.68% 10.28
-A.2.2 B.2 M1-Audio: Audio Spectrogram Transformer (AST)
-We fine-tune the pretrained AST model (MIT/ast-finetuned-audioset), which treats spectrograms as images using Vision Transformer architecture.
-Patch Embedding. The spectrogram (1024128) is divided into 1616 patches:
-Two-Stage Training.
-1.
-Stage 1: Freeze backbone, train classifier (LR=5e-4, 10 epochs)
-2.
-Stage 2: Unfreeze all, fine-tune (LR=5e-6, 15 epochs)
-Limitation: Flattening patches loses the distinction between temporal (rhythm) and frequency (pitch) axes.
-Result: 58.06% 10.54
-A.2.3 B.3 M1-Vision: Vision Transformer (ViT)
-We use a ViT pretrained on facial emotion recognition (dima806/facial_emotions_image_detection).
-Frame-Level Processing. Each of 25 video frames is processed independently:
-Video-Level Prediction. Frame predictions are averaged:
-Limitation: No temporal modeling across frames—cannot capture facial dynamics.
-Result: 75.30% 10.45 (exceeds paper baseline by +0.80pp)
-A.3 C. Model 2: Factorized Attention
-M2 represents our main architectural contribution: axis-specific factorized attention tailored to each modality’s unique structure.
-A.3.1 C.1 Motivation: Why Factorized Attention?
-Standard transformers flatten multi-dimensional data into 1D sequences, losing structural information. For emotion recognition:
-EEG: Spatial (brain regions) vs temporal (dynamics) vs asymmetry (hemisphere differences)
-Audio: Temporal (rhythm) vs frequency (pitch)
-Video: Spatial (face regions) vs temporal (expression dynamics)
-Factorized attention models these axes separately, providing better inductive bias and computational efficiency.
-Complexity Analysis. For video with frames and patches:
-For , :
-A.3.2 C.2 M2-EEG v1: Tri-Stream Spatial-Temporal-Asymmetry Transformer
-Novel Contribution: We introduce a third attention stream for hemispheric asymmetry, based on Davidson’s frontal asymmetry model.
-Architecture. Input EEG (30 channels, 500 samples) passes through convolutional feature extraction, then branches into three parallel attention streams:
-Spatial Stream. Channels attend to channels at each time point:
-This captures which brain regions interact (functional connectivity).
-Temporal Stream. Time points attend to time points for each channel:
-This captures event-related dynamics.
-Asymmetry Stream. We compute difference signals for six hemisphere pairs:
-The F3-F4 pair receives a subtle 1.2 attention weight boost based on its established importance for emotional valence.
-Fusion. Streams are combined via learned softmax weights:
-Skip Connection. A learnable skip preserves convolutional features:
-where with initialized to yield .
-Model Variants:
-| Size | Params | Layers | Heads |
-| Small (recommended) | 138,401 | 2 | 2 |
-| Base | 419,201 | 3 | 4 |
-| Large | 1,626,305 | 4 | 8 |
-Result: 48% (fails—too complex for 280 samples)
-Failure Analysis: With 138K parameters and 280 training samples, the model has 50 more parameters than data points, leading to severe overfitting.
-The sophisticated tri-stream architecture cannot learn meaningful patterns from such limited data.
-A.3.3 C.3 M2-EEG v2: EEGNet-Improved
-As a literature-based alternative, we enhanced EEGNet with proven techniques:
-Squeeze-Excitation Blocks. Learn channel importance:
-Multi-Scale Temporal Convolution. Parallel convolutions with kernels [16, 32, 64] capture different temporal scales.
-Residual Connections. Help gradient flow in deeper networks.
-Additional Improvements: Label smoothing (0.1), AdamW with weight decay (1e-3), cosine annealing LR, data augmentation (amplitude scaling, Gaussian noise, time shift).
-Result: 58.47% 12.56 (approaches paper baseline)
-A.3.4 C.4 M2-Audio: Temporal-Frequency Dual Attention
-Core Idea. Preserve AST’s pretrained features while adding axis-specific attention.
-Grid Reconstruction. Reshape AST’s 1212 patch embeddings into a 10112 grid (time frequency):
-2D Positional Encoding. Add learnable temporal and frequency position embeddings:
-Dual Attention Blocks. Two blocks apply sequential attention:
-Temporal Attention: Each frequency band attends across time:
-Frequency Attention: Each time step attends across frequency:
-Skip Connection. Critical for preserving AST features:
-where initially (88% AST, 12% attention).
-Result: 49.46% (fails—7 subjects stuck at 20% random chance)
-Failure Analysis: Despite preserving 88% of AST features, the additional attention layers introduce noise that corrupts the pretrained representations.
-On small data, even minor corruption cannot be recovered.
-A.3.5 C.5 M2-Vision: Factorized Space-Time Attention
-Architecture. Following ViViT Model 2 (factorized encoder), we apply:
-Spatial Attention. Within each frame, patches attend to patches:
-Temporal Attention. For each patch position, frames attend to frames:
-Skip Connection. Preserves ViT features with learnable .
-Result: 69.54% (best M2 result, but still -5pp below baseline)
-Partial Success Analysis: Skip connection helps preserve pretrained features.
-However, the additional spatial-temporal attention layers still degrade performance compared to simple frame averaging, suggesting that temporal modeling is unnecessary for this dataset’s short video clips.
-A.4 D. Model 3: CNN Improvements
-After M2’s failures, we investigated whether simple fixes to baseline CNNs could outperform complex attention mechanisms.
-A.4.1 D.1 Philosophy: Simple Fixes Over Complex Architectures
-M2’s failures revealed that:
-1.
-Complex models overfit on small datasets
-2.
-Pretrained features are destroyed by additional layers
-3.
-The baseline authors likely tried complex approaches first
-M3 philosophy: Fix bugs, add minimal domain-appropriate features, preserve working architecture.
-A.4.2 D.2 M3-Audio: Delta MFCCs
-Bug Fix: Replace Adam with AdamW (proper weight decay).
-Minimal Addition: Add 40 delta MFCC features (180 220 total):
-Why Not Temporal Features?
-Keeping full temporal MFCCs (40100 = 4000 values) would cause massive overfitting with 280 samples.
-Time-averaged delta features capture dynamics without dimensional explosion.
-Training: AdamW (weight decay 1e-4), label smoothing (0.1).
-Result: 65.56% 9.30 (+3.66pp over paper baseline)
-A.4.3 D.3 M3-EEG v1: Bug Fixes Only
-Critical Bug Found: The original EEGNet classifier contained nn.Softmax(dim=1) before the loss function. Combined with CrossEntropyLoss (which applies log-softmax internally), this computes:
-This double softmax severely damages gradients—the outer softmax on already-normalized probabilities produces near-uniform distributions, destroying learning signal.
-Fix: Remove nn.Softmax from classifier, use raw logits with CrossEntropyLoss.
-Additional Changes: AdamW (weight decay 0.01), label smoothing (0.1).
-Result: 57.66% 13.66 (-2.34pp below paper baseline)
-Analysis: Bug fix alone insufficient. Raw time-domain EEG has poor signal-to-noise ratio (80% artifacts from muscle, eye, and electrical noise).
-A.4.4 D.4 M3-EEG v2: Band Power Features (Best EEG Result)
-Key Insight: Instead of learning frequency decomposition from raw EEG (15,000 values), extract established frequency-domain features first.
-Feature Extraction Pipeline:
-1. Band Power (150 features). For each channel and frequency band , compute power spectral density using Welch’s method:
-where is estimated via:
-Features: 30 channels 5 bands = 150.
-2. Differential Entropy (150 features). For each band-filtered signal assumed Gaussian:
-This is standard in SEED dataset literature. Features: 30 5 = 150.
-3. Alpha Asymmetry (6 features). Following Davidson’s model:
-for pairs (F3,F4), (Fp1,Fp2), (F7,F8), (C3,C4), (P3,P4), (O1,O2).
-Total: 306 features.
-Classifier: Simple MLP:
-Why This Works:
-| Aspect | Raw EEG (v1) | Band Power (v2) |
-| Input dimension | 15,000 | 306 |
-| Signal-to-noise | 20% | High (averaged) |
-| Frequency info | Must learn | Pre-extracted |
-| Asymmetry info | Not explicit | 6 dedicated features |
-Result: 67.62% 7.89 (+7.62pp over paper baseline)
-A.4.5 D.5 M3-Vision v1: SE Block Fix
-Bug Found: The original squeeze-excitation block used reduction ratio :
-With , there is no bottleneck—the “squeeze” does not squeeze! The SE block degenerates to a pointless transformation.
-Fix: Set for proper bottleneck:
-For : (proper squeeze-excitation).
-Result: 69.42% 10.81 (-1.98pp vs paper baseline)
-A.4.6 D.6 M3-Vision v2: Delta Features (Best CNN)
-Following audio success, we add frame-to-frame delta features capturing expression dynamics. The delta features compute the difference between consecutive frame representations:
-where is the frame feature vector from the backbone.
-The final representation combines averaged static and delta features:
-Result: 72.68% 9.56 (+1.28pp over paper CNN baseline)
-A.5 E. Optimization and Regularization
-A.5.1 E.1 Loss Functions
-Cross-Entropy Loss:
-Label Smoothing. Softens targets to prevent overconfidence:
-With and :
-A.5.2 E.2 AdamW Optimizer
-Key difference from Adam: weight decay is applied directly, not scaled by adaptive learning rate.
-A.5.3 E.3 Regularization
-Dropout: Randomly zeros elements with probability :
-Batch Normalization:
-Gradient Clipping: Prevents exploding gradients:
-A.5.4 E.4 Skip Connections in M2
-Critical for preserving pretrained features:
-Initialization: with yields (88% pretrained, 12% new).
-Observation: After training, remains near 0.12—the model learns that pretrained features dominate on small data.
-A.6 F. Model Complexity Analysis
-| Model | Params | Trainable | Result |
-| EEG Models | |||
-| EEGNet (Paper) | 2,600 | 2,600 | 60.0% |
-| M3-EEG v1 (Bug Fix) | 2,600 | 2,600 | 57.66% |
-| M3-EEG v2 (Band Power) | 48K | 48K | 67.62% |
-| M2-EEGNet-Improved | 36,501 | 36,501 | 58.47% |
-| M2-Tri-Stream (small) | 138,401 | 138,401 | 48% |
-| Audio Models | |||
-| AST (M1) | 87M | 4K | 58.06% |
-| M3-Audio (Delta) | 87M | 5K | 65.56% |
-| M2-Dual Attention | 87M | 12M | 49.46% |
-| Vision Models | |||
-| ViT (M1) | 86M | 4K | 75.30% |
-| M3-Vision v1 (SE Fix) | 23M | 23M | 69.42% |
-| M3-Vision v2 (Delta) | 23M | 23M | 72.68% |
-| M2-Factorized | 86M | 12M | 69.54% |
-Rule of Thumb: Trainable parameters should be 10 training samples. With 280 samples, this suggests 2,800 parameters for optimal generalization.
-Observation: M3-EEG v2 succeeds despite 48K parameters because it operates on meaningful 306-dimensional features rather than noisy 15,000-dimensional raw signals.
-A.7 G. Data Augmentation
-EEG:
-Amplitude scaling: ,
-Gaussian noise: ,
-Time shift: Circular shift by 10%
-Audio (SpecAugment):
-Time masking: Zero consecutive frames,
-Frequency masking: Zero consecutive bins,
-Video:
-Horizontal flip with
-Color jitter (brightness, contrast, saturation)
-Random erasing (simulates occlusion)
-A.8 H. Evaluation Metrics
-A.8.1 H.1 Accuracy
-A.8.2 H.2 F1 Score (Weighted)
-Per-class metrics:
-F1 Score:
-Weighted Average:
-where is the number of samples in class .
-A.9 I. Analysis: Why Methods Failed or Succeeded
-A.9.1 I.1 Why M2 Factorized Attention Failed
-1.
-Destruction of Pretrained Features: AST and ViT backbones contain valuable features learned from millions of samples.
-Our attention layers, despite skip connections (), still corrupt these representations.
-2.
-Insufficient Training Data: With 280 samples and 138K+ parameters, M2 models have 50 more parameters than data points. Standard heuristics suggest 10 for good generalization.
-3.
-Complex Inductive Biases Require Data: Factorized attention implements sophisticated biases (separate spatial/temporal reasoning), but learning to use them requires substantial data.
-Evidence: M2 Audio had 7/42 subjects stuck at exactly 20% (random chance for 5 classes), indicating complete failure to learn.
-A.9.2 I.2 Why M3 Simple Fixes Succeeded
-1.
-Domain Knowledge Over Learned Features: Band power features encode decades of neuroscience research into 306 meaningful dimensions. Delta MFCCs encode known speech dynamics.
-2.
-Bug Fixes Restore Intended Behavior: The softmax bug and SE ratio bug were non-obvious errors degrading baseline performance.
-3.
-Minimal Complexity: M3 adds minimal parameters. The 220-feature audio CNN adds only 40 dimensions over baseline.
-Key Insight: With 280 training samples, fixing bugs is more valuable than adding features.